A 5-year Treasury bond has a 3.75% yield. A 10-year Treasury bond yields 6.15%, and a 10-year corporate bond yields 8.55%. The market expects that inflation will average 3.9% over the next 10 years (IP10 = 3.9%). Assume that there is no maturity risk premium (MRP = 0) and that the annual real risk-free rate, r*, will remain constant over the next 10 years. (Hint: Remember that the default risk premium and the liquidity premium are zero for Treasury securities: DRP = LP = 0.) A 5-year corporate bond has the same default risk premium and liquidity premium as the 10-year corporate bond described. The data has been collected in the Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the question below.

What is the yield on this 5-year corporate bond? Round your answer to two decimal places.
Answer #1

DRP+LP = 10-year corporate bond-10-year Treasury bond yield = 6.15-3.75

=2.4%

5-year corporate bond yield = 5-year Treasury bond yield+DRP+LP = 3.75+2.4=6.15%